# Propensity Score Alignment of Unpaired Multimodal Data

### v1

@@ 3 Propensity Score Multi-modal Matching
This says that, even if $\pi_{i}^{(1)}\neq\pi_{i}^{(2)}$, that their relative orderings will still coincide. Then, exact OT will still recover the ground truth matching. See Figure 2 for a visual example of this type of monotonicity. For example, suppose that $t$ is a chemical perturbation of a cell, and thus $\pi_{i}^{(1)}$, $\pi_{j}^{(2)}$ can be seen as a measure of biological response to the perturbation, e.g., in a treated population, $\pi_{i_{1}}>\pi_{i_{2}}$ indicates samples $i_{1}$ had a stronger response than sample ${i_{2}}$, as perceived by the first modality indexed by $i$. Then, this monotonocity states that we should see the same $\pi_{j_{1}}>\pi_{j_{2}}$ in the other modality as well, if the samples $i_{1}$ and $i_{2}$ truly corresponded to $j_{1}$ and $j_{2}$.
When $t$ is not a binary treatment and the propensity scores are multidimensional, a condition known as cyclic monotonicity is sufficient for recovering the true matching, see Appendix A for details. However, unlike the 1-d case, we are not aware of more intuitive conditions under which we should expect cyclic monotonicity to hold.
4 Methodology
For the remainder, we will drop $(e,t)$ from our notation, and use $(x_{i},t_{i})$ to denote observations from modality 1, and $(x_{j},t_{j})$ to denote observations from modality 2 whenever it does not cause confusion to do so. Given a multimodal dataset with observations $\{(x_{i},t_{i})\}_{i=1}^{n_{1}}$ and $\{(x_{j},t_{j})\}_{j=1}^{n_{2}}$, we wish to compute a matching matrix (or coupling) between the two modalities. We define a $n_{1}\times n_{2}$ matching matrix $M$ where $M_{ij}$ represents the likelihood of $x_{i}$ being matched to $x_{j}$. We always perform matching only within observations with the same value of $t$, so that in practice we obtain a matrix $M_{t}$ for each $t$.
$\pi_{i}\approx\pi(x_{i})=P(T=t\mid X_{i}^{(e,t)}=x_{i}).$ (9)
This results in the transformed datasets $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$, where $\pi_{i}$, $\pi_{j}$ are in the $T$ dimensional simplex. We use this correspondence to compute a cross-modality distance function:
$d(x_{i},x_{j}):=d^{\prime}(\pi_{i},\pi_{j}),$ (10)
where in practice, we typically compute the Euclidean distance in $\mathbb{R}^{T}$ of the logit-transformed classification scores. Given this distance function, we rely on existing matching techniques to constructing a matching matrix. In our experiments, we found OT matching gave the best performance, but we also evaluated Shared Nearest Neighbour matching; details of the latter can be found in Appendix B.
4.1 Optimal Transport Matching
The propensity score distance allows us to easily compute a cost function associated with transporting mass between modalities, $c(x_{i},x_{j})=d^{\prime}(\pi_{i}\pi_{j})$. Let $p_{1},p_{2}$ denote the uniform distribution over $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$ respectively. Discrete OT aims to solve the problem of optimally redistributing mass from $p_{1}$ to $p_{2}$ in terms of incurring the lowest cost. Let $C_{ij}=c(x_{i},x_{j})$ denote the $n_{1}\times n_{2}$ cost matrix. The Kantorovich formulation of optimal transport aims to solve the following constrained optimization problem:
$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij},\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2}.$
$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij}-\lambda H(M),\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2},$
5 Cross-modality prediction
Given a matching matrix $M$, we can interpret this as the probability that each sample, $i$, from modality (1) is matched to sample $j$ in modality (2); that is $M_{i,j}=P(x^{(2)}_{j}|x^{(1)}_{i})$. As a downstream task, and as an evaluation metric for $M$, we can use this matching to estimate a cross-modal prediction model, $f_{\theta}$, that maps from modality (2) to (1) by minimizing the following loss,
$\mathcal{L}(\theta):=\sum_{i}(x_{i}^{(1)}-M_{i}f_{\theta}(x_{j}^{(2)}))^{2}.$ (11)
However, this requires evaluating $f_{\theta}$ for all $n_{2}$ examples from modality (2) for each of the $n_{1}$ examples in modality (1). Of course, we can avoid this cost with stochastic gradient descent by sampling from modality $(2)$ via $M_{i\cdot}$ for each training example $(1)$, but to get an unbiased estimate of $\nabla_{\theta}\mathcal{L}$, we need two samples from modality (2) for each sample from modality (1),
$\nabla\mathcal{L}(\theta)\approx$ $-2\left(x_{i}^{(1)}-f_{\theta}(\dot{x}_{j}^{(2)})\right)\nabla_{\theta}f_{\theta}(\ddot{x}_{j}^{(2)})\quad$ $\dot{x}_{j}^{(2)},\ddot{x}_{j}^{(2)}\sim P(x^{(2)}_{j}|x^{(1)}_{i}).$ (12)
By taking two samples as in equation (12), we get an unbiased estimator of $\nabla\mathcal{L}(\theta)$, whereas a single sample would have resulted in optimizing an upper-bound on equation (11); for details, see Hartford et al. (2017) where a similar issue arises in the gradient of their causal effect estimator.
6 Experiments
We evaluate our proposed methodology on two datasets. The first is a synthetic interventional image dataset generated satisfying the assumptions of Section 3. The second is a real-world single-cell CITE-seq data obtained from the NeurIPS 2021 Multimodal single-cell data integration competition (Lance et al., 2022), which provides a ground truth matching by allowing for a small number of cell surface proteins to be measured simultaneously to RNA sequencing. Note the CITE-seq dataset is not interventional—we use the cell type as the classification target $t$ instead. All details are made available in the Appendix. In both cases the ground truth matching is known to make evaluation possible, but hidden during training.
Baselines   In terms of the problem setting, the most closely related method is Gromov-Wasserstein optimal transport (SCOT) (Demetci et al., 2022), which explicitly computes a cross-modality cost for optimal transport matching by leveraging the local geometry within each modality. For the CITE-seq data, graph-linked VAE (scGLUE), which has access to specific biological metadata that connects the biological modalities, can also be used for matching. [^2] A variant of scGLUE was the top entrant in the NeurIPS 2021 Multi-modal single-cell data integration competition, from which we source our data. However, ground truth matchings were made available in the training set in that case, which fundamentally changes the nature and difficulty of the problem. However, neither of the preceding approaches use the experimental label information, which is a key component of our method. In fact, the only method to our knowledge that leverages such information is the VAE approach developed in Yang et al. (2021), which also minimizes a classification loss from the latent space, which we will consider our main baseline. For all methods besides SCOT, we use both SNN and OT to perform the final matching in the latent space. Note that when combined with OT, our VAE baseline also resembles Cao et al. (2022).
Model Details   For our method and the VAE, we use the general architecture of a linear classification head on top of a suitable encoder. The architecture of this encoder is always the same between the VAE and our method. For the image dataset, we use the convolutional neural network from Yang et al. (2021) as the encoder, and for the CITE-seq dataset, we use fully-connected MLPs for both RNA (top 200 PC’s) and protein data. Optimization was performed using Adam with a one-cycle learning rate scheduler. We use the standard cross-entropy loss to train our classifiers for propensity score estimation. For other methods, we use existing implementations with suggested default settings.
Matching Details   Both SNN and OT use the Euclidean distance function to determine neighbours and compute the cost matrix, respectively. SCOT uses the correlation distance by default, and we found that this resulted in better performance than Euclidean distance. We use only a single neighbour for SNN matching, which interestingly resulted in the best performance. Both SCOT and OT solve the entropic regularized OT, for these we use a regularization parameter of $0.05$.
Method MSE Trace
(Med (Q1, Q3)) (Med (Q1, Q3)) ×$\times 10^{-3}$
SCOT 0.0354 0.5964
VAE+SNN 0.0622 3.116
(0.0571, 0.0676) (2.818, 3.213)
VAE+OT 0.0324 7.733
(0.0316, 0.0350) (7.473, 7.794)
PS+SNN 0.0552 7.924
(0.0530, 0.0558) (7.569, 9.504)
PS+OT 0.0316 18.329
(0.0300, 0.0330) (17.068, 18.987)
Rand 0.0709 N/A
(0.0707, 0.0714)
Table 1: Results on synthetic image data. Metrics are computed on a held out test set over 12 groups corresponding to interventions on the latent position, with approximately 1700 observations per group. A trace of $10$ for example corresponds to a total weight of $1710\times 10^{-3}\times 1700=$ 17 on the true matching, out of 1700.
Method FOSCTTM Trace
(Median (Q1, Q3)) (Median (Q1, Q3))
SCOT 0.4596 0.0200
GLUE+SNN 0.4412 0.0362
GLUE+OT 0.5309 0.0323
VAE+SNN 0.3816 0.0612
(0.3760, 0.3822) (0.0588, 0.0634)
VAE+OT 0.3953 0.0814
(0.3912, 0.4045) (0.0777, 0.8895)
PS+SNN 0.3126 0.0941
(0.3121, 0.3160) (0.0880, 0.0989)
PS+OT 0.3049 0.1163
(0.3008, 0.3078) (0.1093, 0.1250)
Table 2: Results on CITE-seq data. Metrics are computed on a held out test set averaged over a total of 45 cell types (groups) with a varying amount of observations per group. As such, the average trace can be difficult to interpret (for example, the maximum trace is near $1$, corresponding to a near-perfect matching within a group with only $7$ cells in the test set), but we remark that OT matching on propensity scores outperforms the other methods within each group as well as on average.
Evaluation Metrics   We report three evaluation metrics. The trace metric computes the average mass that the matching matrix places on the true matches (higher is better), and FOSCTTM reports the Fraction Of Samples Closer Than the True Match (FOSCTTM) ((Demetci et al., 2022), (Liu et al., 2019)) (lower is better, 0.5 corresponds to random guessing). For synthetic images where we know the ground true latent, we report the MSE to the true latent after matching. Full details of these metrics are given in Section D.1.
Modality Prediction   We also consider one metric that examines whether matched samples are useful for downstream tasks. For this, we chose the cross-modality prediction task of predicting protein levels from RNA expression in the CITE-seq dataset. To do this, we train a supervised learning model (in our case, a 2-layer MLP with either MSE loss or the unbiased gradients loss in Section 5) using samples generated by different matching matrices $M$. As baselines, we also train on pairs from the ground truth matching ($M_{ii}=1$) and random sampling ($M_{ij}=1/n$). In view of the performance observed in Table 1 and Table 2, we only considered $M$ resulting from matching with OT using the propensity score and VAE embeddings.
Method $R^{2}$ (MSE) $R^{2}$ (Unbiased)333 This refers to training the predictor with the unbiased MSE (Equation 11). The evaluation is still using $R^{2}$ as defined by the standard MSE.
(Med (Q1, Q3)) (Med (Q1, Q3))
Rand 0.1383 0.1727
(0.1372, 0.1402) (0.1701, 0.1731)
VAE+OT 0.1493 0.1142
(0.1179, 0.1724) (0.0786, 0.1594)
PS+OT 0.2174 0.2331
(0.2062, 0.2228) (0.2069, 0.2504)
True Pairs 0.2243 N/A [^4] The standard and unbiased MSE are equivalent in this case.
(0.2234, 0.2257)
Table 3: Modality prediction results on CITE-seq data. Entries represent $R^{2}$ computed on a held-out test set consisting of ground-truth pairs, which is computed as $1-\text{MSE}(\hat{x},x)/\text{MSE}(\bar{x},x)$, where $\hat{x}$ are the predicted protein levels, and $\bar{x}$ is the mean value of the validation set. Note random sampling, especially with the unbiased gradient loss, achieves a positive $R^{2}$ against $\bar{x}$ because it samples randomly within $t$, which results in learning group-level statistics.
6.1 Validation Cross-Entropy Monitor
When the assumptions are satisfied, the success of our propensity score method hinges on how well we can learn the conditional distributions $p(t\mid X_{i}^{(e,t)})$. Thus, the validation cross entropy can be used as a tractable metric that can serve as a model selection monitor for the matching performance, which would be unknown in practice. The effectiveness of such a monitor is validated in our experiments, even on real-world CITE-seq data, where we see that a lower validation loss typically corresponds to higher matching performance which is not the case for the VAE (Figure 3). We suspect that this is precisely due to the additional requirement that the VAE minimize a reconstruction loss, in addition to the classification cross-entropy.
6.2 Results
We checkpoint the classifier and VAEs at the lowest validation loss and report their metrics on a held out test set over 10 random seeds. Note that this does not necessarily select the embedding model that exhibits optimal matching, but instead the best model to select without validating against ground truth matching metrics, see Figure 3. We further used 10 random seeds for training the MLP used for modality prediction. scGLUE is trained according to the public implementation, which uses learning rate reduction and early stopping strategies. SCOT is a non-iterative approach and thus we directly report its results.
We found that OT matching on propensity scores consistently outperforms other methods on all metrics, typically followed by SNN matching on propensity scores, or OT matching on VAE embeddings. Furthermore, both the VAE and propensity score, which leverage experimental label information, tend to perform better. This suggests that using the propensity score as embeddings for matching, and using OT to perform the final matching both independently improve matching performance. Curiously, SCOT performs well on the MSE metric for image data, but only places slightly above random in the trace metric. This indicates that it matches scenes with similar latent coordinates, without placing any significant weight on the exact ground truth. This suggests that exact matches may not be necessary for a method to be useful for downstream tasks.
Surprisingly, using the propensity score embeddings with OT matching appears to improve generalization in the modality prediction task (Table 3), when using the unbiased MSE over a model trained on ground truth pairings (indeed, we observed that the ground truth model had a lower training loss, but higher test loss). This reveals an unexpected benefit of (soft) matching: we can sample from the conditional to minimize the loss Equation 11, which, with a suitable $M$, results in improved generalization compared to the naive MSE. This hinges strongly on the quality of $M$—the $M$ resulting from the VAE embeddings results in worse generalization, and in practice the quality of the matching, at least in absolute terms, remains hidden.
7 Conclusion
This work presents a simple algorithm for aligning unpaired data from different modalities. The method is both very general—only requiring a classifier to be trained on each modality—and highly effective for matching, which we show both theoretically and empirically. As a downstream task, we demonstrate the effectiveness of the resulting matchings for cross modality prediction, which leads to better generalization than the ground truth matching on the dataset we evaluated. We suspect that this improved generalization is the result of implicitly enforcing invariance to modality specific information, but more work is needed to evaluate the conditions under which this improved generalization occurs.
Appendix A Cyclic monotonicity
We can see the monotonicity requirement Equation 8 as the monotonicity of the function with graph $(\pi_{i}^{(1)},\pi_{i}^{(2)})\in[0,1]^{2}$. In higher dimensions, we require that the “graph” satisfies the following cyclic monotonicity property (Villani, 2009):
@@ Appendix A Cyclic monotonicity
Note in our setting, the OT cost function is the Euclidean distance, $c(x,y)=\|x-y\|_{2}$. It is known that the OT solution must satisfy cyclic monotonicity. Thus, if the true pairing is uniquely cyclically monotone, we can recover it with OT.
@@ Appendix B Shared Nearest Neighbours Matching
Using the propensity score distance, we can compute nearest neighbours both within and between the two modalities. We follow Cao & Gao (2022) and compute the normalized shared nearest neighbours (SNN) between each pair of observations as the entry of the matching matrix. For each pair of observations $(\pi_{i}^{(1)},\pi_{j}^{(2)})$, we define four sets:
@@ Appendix C Omitted Proofs
Let $x^{(e,t)}$ denote the observed modality and $z^{(t)},u^{(e)}$ be the unique corresponding latent values. By injectivity,
$\pi(x^{(e,t)})$ $=P(t|X^{(e,t)}=x^{(e,t)})$
$=P(t|Z^{(t)}=z^{(t)},U^{(e)}=u^{(e)})$
$=P(t|Z^{(t)}=z^{(t)})=\pi(z^{(t)}),$ (17)
for $e=1,2$, since we assumed $U^{(e)}{\perp\!\!\!\perp}t\mid Z^{(t)}$ in Section 3. Since this holds pointwise, it shows that $\pi(X^{(1,t})=\pi(X^{(2,t)})=\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables. Now, a classical result of Rubin (1974) gives that $Z^{(t)}{\perp\!\!\!\perp}t\mid\pi(Z^{(t)})$, and that for any other function $b$ (a balancing score) such that $Z^{(t)}{\perp\!\!\!\perp}t\mid b(Z^{(t)})$, we have $\pi(Z^{(t)})=g(b(Z^{(t)}))$. The first property written in information theoretic terms yields,
$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (18)
since $\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables, as required. ∎
@@
where $C=\log(p(t=i))-\log(p(t=0))$ is constant in $z$, and that $h(\pi(z))_{0}\equiv 0$. Ignoring the constant first dimension, we can view $h$ as an invertible map to $\mathbb{R}^{T}$. Under this convention, the map $h\circ\pi:\mathbb{R}^{d}\to\mathbb{R}^{T}$ is smooth ($\log$ is smooth, and the densities are smooth by assumption). Since it is smooth, it cannot be injective if $T<d$ (Spivak, 2018). Finally, since $h$ is bijective, this implies that $\pi$ cannot be injective. ∎
@@ D.1 Evaluation Metrics
We can evaluate how well samples are matched using the ground truth provided by our datasets. In these cases, the dataset sizes are necessarily balanced, so that $n=n_{1}=n_{2}$. In each case, the metric is a function of a $n\times n$ matching matrix $M$, which recall we compute within samples with the same $t$. Our reported results are then averaged over each cluster. We randomize the order of the datasets before performing the matching to avoid pathological cases.
Trace Metric   Assuming the given indices correspond to the true matching, we can compute the average weight on correct matches, which is the normalized trace of $M$:
@@
Latent MSE   Because the image dataset is synthetic, we have access to the ground truth latent values that generated the images, $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$. We compute matched latents as $M\mathbf{z}$, the mean projection according to the matching matrix. Then, to evaluate the quality of the matching, we compute the MSE:
@@
FOSCTTM   For the CITE-seq dataset, we can use the Fraction Of Samples Closer Than the True Match (FOSCTTM) (Demetci et al., 2022; Liu et al., 2019) as an alternative matching metric. First, we distribute the mass of $\mathbf{x}^{(2)}=\{x_{j}\}_{j=1}^{n}$ to $\mathbf{x}^{(1)}=\{x_{i}\}_{i=1}^{n}$ as $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$, resulting in a projection of the first modality $\mathbf{x}^{(1)}$ in the space of the second modality. Then, we can compute a cross-modality distance as follows. For each point in $\hat{\mathbf{x}}^{(1)}$, we compute the Euclidean distance to each point in $\mathbf{x}^{(1)}$, and compute the fraction of samples in $\mathbf{x}^{(1)}$ that are closer than the true match. We also repeat this for each point in $\mathbf{x}^{(1)}$, computing the fraction of samples in $\hat{\mathbf{x}}^{(1)}$ in this case. That is, assuming again that the given indices correspond to the true matching, we compute:
@@
where notice that this is a function of $M$ through the computation $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$. As a baseline, we should expect a random matching, when distances between points are randomly distributed, to have an FOSCTTM of $0.5$.

### v2

@@ Appendix A Relaxing (A1)
This says that, even if $\pi_{i}^{(1)}\neq\pi_{i}^{(2)}$, that their relative orderings will still coincide. Then, exact OT will still recover the ground truth matching. See Fig. 3 for a visual example of this type of monotonicity. For example, suppose that $t$ is a chemical perturbation of a cell, and thus $\pi_{i}^{(1)}$, $\pi_{j}^{(2)}$ can be seen as a measure of biological response to the perturbation, e.g., in a treated population, $\pi_{i_{1}}>\pi_{i_{2}}$ indicates samples $i_{1}$ had a stronger response than sample ${i_{2}}$, as perceived by the first modality indexed by $i$. Then, this monotonocity states that we should see the same $\pi_{j_{1}}>\pi_{j_{2}}$ in the other modality as well, if the samples $i_{1}$ and $i_{2}$ truly corresponded to $j_{1}$ and $j_{2}$.
A.1 Cyclic Monotonicity
We can see the monotonicity requirement Eq. 14 as the monotonicity of the function with graph $(\pi_{i}^{(1)},\pi_{i}^{(2)})\in[0,1]^{2}$. In higher dimensions, we require that the “graph” satisfies the following cyclic monotonicity property [Villani, 2009]:
@@ A.1 Cyclic Monotonicity
Note in our setting, the OT cost function is the Euclidean distance, $c(x,y)=\|x-y\|_{2}$. It is known that the OT solution must satisfy cyclic monotonicity. Thus, if the true pairing is uniquely cyclically monotone, we can recover it with OT. However, we are unaware of common violations of (A1) that would satisfy cyclic monotonicity.
@@ Appendix B Shared Nearest Neighbours Matching
Using the propensity score distance, we can compute nearest neighbours both within and between the two modalities. We follow Cao and Gao [2022] and compute the normalized shared nearest neighbours (SNN) between each pair of observations as the entry of the matching matrix. For each pair of observations $(\pi_{i}^{(1)},\pi_{j}^{(2)})$, we define four sets:
@@ Appendix C Proofs
Let $x^{(e)}$ denote the observed modality and $z,u^{(e)}$ be the unique corresponding latent values. By injectivity,
$\pi(x^{(e)})$ $=P(t|X^{(e)}=x^{(e)})$
$=P(t|Z=z,U^{(e)}=u^{(e)})$
$=P(t|Z=z)=\pi(z),$ (19)
for $e=1,2$, since we assumed $U^{(e)}{\perp\!\!\!\perp}t\mid Z$. Since this holds pointwise, it shows that $\pi(X^{(1})=\pi(X^{(2)})=\pi(Z)$ as random variables. Now, a classical result of Rubin [1974] gives that $Z{\perp\!\!\!\perp}t\mid\pi(Z)$, and that for any other function $b$ (a balancing score) such that $Z{\perp\!\!\!\perp}t\mid b(Z)$, we have $\pi(Z)=g(b(Z))$. The first property written in information theoretic terms yields,
$I(t,Z\mid\pi(Z))=I(t,Z\mid\pi(X^{(e)}))=0,$ (20)
since $\pi(X^{(e)})=\pi(Z^{(t)})$ as random variables, as required. ∎
@@
where $C=\log(p(t=i))-\log(p(t=0))$ is constant in $z$, and that $h(\pi(z))_{0}\equiv 0$. Ignoring the constant first dimension, we can view $h$ as an invertible map to $\mathbb{R}^{T}$. Under this convention, the map $h\circ\pi:\mathbb{R}^{d}\to\mathbb{R}^{T}$ is smooth ($\log$ is smooth, and the densities are smooth by assumption). Since it is smooth, it cannot be injective if $T<d$ [Spivak, 2018]. Finally, since $h$ is bijective, this implies that $\pi$ cannot be injective. ∎
@@ D.1 Evaluation Metrics
D.1.1 Known Ground Truth
In the synthetic image and CITE-seq datasets, a ground truth matching is known, and we can evaluate the quality of the synthetic matching directly against the truth. In these cases, the dataset sizes are necessarily balanced, so that $n=n_{1}=n_{2}$. In each case, we evaluate the quality of our $n\times n$ matching matrix $M$, which we compute within samples with the same $t$. Our reported results are then averaged over each cluster. Note we randomize the order of the datasets before performing the matching to avoid pathologies.
Trace Metric   Assuming the sample indices correspond to the true matching, we can compute the average weight on correct matches, which is the normalized trace of $M$:
@@ D.1.1 Known Ground Truth
Latent MSE   On the image dataset, we have access to the ground truth latent values that generated the images, $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$. We compute matched latents as $M\mathbf{z}$, the barycentric projection according to the matching matrix. Then, to evaluate the quality of the matching in terms of finding similar latents, we compute the MSE:
@@
FOSCTTM   We do not have access to ground truth latents in the CITE-seq dataset, so use the Fraction Of Samples Closer Than the True Match (FOSCTTM) [Demetci et al., 2022, Liu et al., 2019] as an alternative matching metric. First, we use $M$ to project $\mathbf{x}^{(2)}=\{x_{j}\}_{j=1}^{n}$ to $\mathbf{x}^{(1)}=\{x_{i}\}_{i=1}^{n}$ as $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$. Then, we can compute a cross-modality distance as follows. For each point in $\hat{\mathbf{x}}^{(1)}$, we compute the Euclidean distance to each point in $\mathbf{x}^{(1)}$, and compute the fraction of samples in $\mathbf{x}^{(1)}$ that are closer than the true match. We also repeat this for each point in $\mathbf{x}^{(1)}$, computing the fraction of samples in $\hat{\mathbf{x}}^{(1)}$ in this case. That is, assuming again that the given indices correspond to the true matching, we compute:
@@
where notice that this evaluates $M$ through the computation $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$. As a baseline, we should expect a random matching, when distances between points are randomly distributed, to have an FOSCTTM of $0.5$.
Prediction Accuracy   We also trained a cross-modality prediction (translation) model $f_{\theta,M}$ to predict CITE-seq protein levels from gene expression based on matched pseudosamples. Let $\mathbf{x^{(1)}}=\{x_{i}\}$, $\mathbf{x^{(2)}}=\{x_{j}\}$ denote protein and gene expression, respectively. We trained a simple 2-layer MLP minimizing either the standard MSE, using pairs $(x_{i},\hat{x}_{j})$, $\hat{x}_{j}\sim M_{i\cdot}$, or following the projected loss with unbiased estimates in Section 5. Each batch in general consists of samples from all $t$, but the $\hat{x}_{j}$ sampling step occurs within the perturbation. Let $\hat{\mathbf{x}}^{(1)}_{test}=\{f_{\theta,M}(x_{j})\}$. We report the $R^{2}$ on a randomly held-out test set of ground truth pairs (again, consisting of samples from all $t$), which is defined as the following:
$R^{2}(f_{\theta,M})=\frac{MSE(\mathbf{x}^{(1)}_{test},\hat{\mathbf{x}}^{(1)}_{test})}{MSE(\mathbf{x}^{(1)}_{test},\bar{\mathbf{x}}^{(1)}_{test})},$ (28)
where $\bar{\mathbf{x}}^{(1)}_{test}$ is the naive mean (over all perturbations) estimator which acts as a baseline.
D.1.2 Unknown Ground Truth
We train a cross-modality prediction model to predict gene expression from cell images based on matched pseudosamples in the same way as in CITE-seq, but only using the projected loss with unbiased estimates. Denote this model for a matching matrix $M$ by $f_{\theta,M}$.
Because we do not have access to ground truth pairs within each perturbation, we resort to distributional metrics. Let $\mathbf{x^{(1)}}_{t}=\{x_{i,t}\}_{i=1}^{n_{t1}}$, $\mathbf{x^{(2)}}_{t}=\{x_{j,t}\}_{j=1}^{n_{t2}}$ denote gene expression and cell images in a held out test set respectively in perturbation $t$. Let $\hat{\mathbf{x}}^{(1)}_{t}=\{f_{\theta,M}(x_{j,t})\}_{j=1}^{n_{t2}}$. We compute empirical versions of statistical divergences
$D_{t}(f_{\theta,M}):=D(\mathbf{x}^{(1)}_{t},\hat{\mathbf{x}}^{(1)}_{t}),$ (29)
where $D$ is either the L2 norm of the difference in empirical mean, empirical Kullback-Leibler divergence or 1-Wasserstein distance. We report these weighted averages of $D_{t}$ over the perturbations $t$ according to the number of samples in the modality of prediction interest.